OpenStudy (anonymous):

find the anti derivative (1-x)(x-4)dx

OpenStudy (anonymous):

multiply out first then use the power rule backwards

OpenStudy (anonymous):

so it's \[-\frac{ 1 }{ 3 }x ^{3}+\frac{ 5 }{ 2 }x ^{2}-2x\] correct?

OpenStudy (dumbcow):

(1-x)(x-4) = -x^2 +5x -4 anti-derivative is: -(1/3)x^3 +(5/2)x^2 -4x + C

OpenStudy (anonymous):

oh i see where i went wrong. did something stupid lol

OpenStudy (anonymous):

i see you agree lol

OpenStudy (dumbcow):

yep :)